### v2

@@ Abstract.
We present two further applications of symplectic group symmetry. First, we show that the PPT-squared conjecture holds within the class of PPT linear maps that are either symplectic-covariant or conjugate-symplectic-covariant. Second, we resolve a conjecture of Pál and Vértesi concerning the optimal lower bound of the Sindici–Piani semidefinite program for PPT entanglement.
@@ 1. Introduction
In the case of the $d\otimes d$ system, evidence for the existence of PPT states whose Schmidt number scales linearly in $d$ was first obtained in [SWZ11] using tools from asymptotic geometric analysis, although explicit constructions remained elusive. Early explicit examples exhibited PPT states whose Schmidt number grows logarithmically in $d$ [CYT17], while subsequent works improved this bound to linear scaling in $d$ [HLLMH18, PV19, Car20]. At present, the best known general upper bound is $\lfloor d/2\rfloor$, except for the dimensions $d=3$ [Cho82, Hor97] and $d=5$ [KG24], and explicit constructions achieving Schmidt number of order $d/2$ has been known only in limited settings [PV19, Car20]. The well-known dual formulation of this problem asks for the existence of $k$-positive indecomposable (i.e., not of the form $\mathcal{L}^{\prime}+\top\circ\mathcal{L}^{\prime\prime}$ for some completely positive maps $\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime\prime}$) linear maps on $M_{d}(\mathbb{C})$ as large as possible, a question originally raised in [Ter01]. This dual viewpoint has also driven much of the development of the theory of $k$-positive maps and its Choi matrices [CK09, YLT16, HLLMH18, RBO20, MOM25, HK25a, HK25b, vEKD25].
@@
In this work, we contribute to both of the above questions on $k$-positivity and Schmidt numbers by exploring new classes of quantum objects possessing symmetry with respect to the symplectic group. More precisely, we study linear maps that are covariant under conjugation by unitary symplectic matrices $S$, as well as bipartite quantum states that are invariant under $S\otimes S$ or $S\otimes\overline{S}$ actions, each parametrized by two real variables. The representation theory of the symplectic group gives rise to algebraic structures that are formally similar to those arising in symmetry classes associated with the orthogonal group [VW01, PY24]; however, the resulting entanglement properties turn out to be drastically different. As we show, this symmetry setting supports a remarkably rich entanglement landscape, including highly entangled PPT states and highly positive yet indecomposable linear maps.
@@
In particular, we emphasize a new construction of Breuer–Hall–type maps, which are shown to be $k$-positive and indecomposable in strong sense and thus remedy a limitation of the original Breuer–Hall map.
@@
In this sense, we may refer to $\mathcal{L}_{k}^{\rm BH}$ as the $k$-Breuer–Hall maps. We remark that the construction of optimal $k$-positive indecomposable maps has remained largely out of reach beyond the case $k=1$; here, however, the group symmetry method leads to a substantial simplification of the argument. Furthermore, we can provide general Schmidt number criteria (Proposition 4.6) and additional examples of PPT states with high Schmidt number (Corollaries 4.8 and 4.9) from these linear maps.
@@ 2.2. Compact group symmetries and twirling operations
2.3. Symplectic group symmetry
@@ 3.2. $k$-Positivity
[…]
@@ 6. Optimal lower bounds for the Sindici–Piani semidefinite program
by the assumption of $\sigma$. This shows the relation (45). In particular, this implies that the optimal state $\sigma^{*}$ for $p^{\rm PPT}(|\omega_{d}\rangle\langle\omega_{d}^{\Omega}|)$ can be chosen to be of the form
@@
Acknowledgements: The author thanks Aabhas Gulati and Seung-Hyeok Kye for the helpful discussions and comments. We would also like to thank the organizers of the first meeting of Operator Algebras and Quantum Information for the kind invitation to present a preliminary version of this work. S.-J.P was supported by the ANR project ESQuisses, grant number ANR-20-CE47-0014-01.

### v3

@@ Abstract.
We present two further applications of symplectic group symmetry. First, we show that the PPT-squared conjecture holds within the class of PPT linear maps that are either symplectic covariant or conjugate symplectic covariant. Second, we resolve a conjecture of Pál and Vértesi concerning the optimal lower bound of the Sindici–Piani semidefinite program for PPT entanglement.
@@ 1. Introduction
In the case of the $d\otimes d$ system, evidence for the existence of PPT states whose Schmidt number scales linearly in $d$ was first obtained in [SWZ11] using tools from asymptotic geometric analysis, although explicit constructions remained elusive. Early explicit examples exhibited PPT states whose Schmidt number grows logarithmically in $d$ [CYT17], while subsequent works improved this bound to linear scaling in $d$ [HLLMH18, PV19, Car20]. At present, the best known general upper bound is $\lfloor d/2\rfloor$, except for the dimensions $d=3,5,7$ and $9$ [Cho82, Hor97, KG24, KG25], and explicit constructions achieving Schmidt number of arbitrary order $d/2$ has been known only in limited settings [PV19, Car20]. The well-known dual formulation of this problem asks for the existence of $k$-positive indecomposable (i.e., not of the form $\mathcal{L}^{\prime}+\top\circ\mathcal{L}^{\prime\prime}$ for some completely positive maps $\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime\prime}$) linear maps on $M_{d}(\mathbb{C})$ as large as possible, a question originally raised in [Ter01]. This dual viewpoint has also driven much of the development of the theory of $k$-positive maps and its Choi matrices [CK09, YLT16, HLLMH18, RBO20, MOM25, HK25a, HK25b, vEKD25].
@@
In this work, we contribute to both of the above questions on $k$-positivity and Schmidt numbers by exploring new classes of quantum objects possessing symmetries with respect to the symplectic group. More precisely, we study linear maps that are covariant under conjugation by unitary symplectic matrices $S$, as well as bipartite quantum states that are invariant under $S\otimes S$ or $S\otimes\overline{S}$ actions, each parametrized by two real variables. The representation theory of the symplectic group gives rise to algebraic structures that are formally similar to those arising in symmetry classes associated with the orthogonal group [VW01, PY24]; however, the resulting entanglement properties turn out to be drastically different. As we show, this symmetry setting supports a remarkably rich entanglement landscape, including highly entangled PPT states and highly positive yet indecomposable linear maps.
@@
In particular, we emphasize a new construction of Breuer–Hall–type maps, which are shown to be $k$-positive and indecomposable in stronger sense, and thus remedy a limitation of the original Breuer–Hall map.
@@
In this sense, we may refer to $\mathcal{L}_{k}^{\rm BH}$ as the $k$-Breuer–Hall maps. We remark that the construction of $k$-positive linear maps possessing either of the above two properties has remained largely out of reach beyond the case $k=1$; here, however, the group-symmetry method leads to a substantial simplification of the argument. Furthermore, from these linear maps we derive general Schmidt number criteria (Proposition 4.6) and obtain additional examples of PPT states with high Schmidt number (Corollaries 4.8 and 4.9).
@@ 2.2. Compact group symmetries and twirling operations
2.3. Symplectic group symmetries
@@ 3.2. $k$-Positivity
    For $k=1,\ldots,d/2-1$, the map $\mathcal{L}_{p,q}$ is $k$-positive but cannot be written in the form $\mathcal{L}^{\prime}+\top\circ\mathcal{L}^{\prime\prime}$ for any $(k+1)$-positive linear maps $\mathcal{L}^{\prime}$ and $\mathcal{L}^{\prime\prime}$ if and only if
@@ 6. Optimal lower bounds for the Sindici–Piani semidefinite program
by the assumption of $\sigma$. This shows the relation (45). In particular, this implies that the optimal state $\sigma^{*}$ for $p^{\rm PPT}(|\omega_{d}^{\Omega}\rangle\langle\omega_{d}^{\Omega}|)$ can be chosen to be of the form
@@
7. Conclusion
By exploiting symplectic group symmetries, we construct a broad class of new $k$-positive indecomposable linear maps and PPT states exhibiting a high degree of entanglement, attaining the best known general bounds from [PV19, Car20]. In particular, we provide a systematic construction of nontrivial quantum objects, including $k$-positive maps with strong indecomposability properties and PPT entangled states displaying an extremal Schmidt number gap between a state and its partial transpose, a phenomenon that had remained largely beyond reach prior to this work. Furthermore, we demonstrate several applications of our framework, including results related to the PPT-squared conjecture and the optimal bound in the SDP associated with PPT entanglement.
Our results demonstrate the effectiveness of the symmetry method as a systematic tool for analyzing nontrivial phenomena related to Schmidt numbers and positivity. By exploiting symplectic group symmetries and tools from projective geometry, we obtain analytically tractable families in which representation-theoretic structures translate directly into geometric descriptions of entanglement, revealing an unexpected interplay between representation theory, projective geometry, and quantum information. Remarkably, even a simple two-parameter family already yields a broad class of highly entangled PPT states and highly positive yet indecomposable maps, suggesting that symplectic symmetry captures essential features of extremal entanglement behavior. This observation naturally motivates further exploration beyond the unitary symplectic group, for instance toward weaker symmetry assumptions such as orthogonal symplectic symmetry or more general covariance classes. We expect that such extensions will uncover new entanglement structures and provide a unified framework for constructing and classifying quantum states, potentially improving the current best bounds and further leading to counterexamples to several important open problems, such as the PPT-squared conjecture [Chr12, CMHW19] and the NPT distillation problem [HRŻ22]. We leave these directions for future work.
Acknowledgements: The author thanks Aabhas Gulati, Seung-Hyeok Kye, Robin Krebs, and Mariami Gachechiladze for the helpful discussions and comments. We would also like to thank the organizers of the first meeting of Operator Algebras and Quantum Information for the kind invitation to present a preliminary version of this work. S.-J.P. was supported by the ANR project ESQuisses, grant number ANR-20-CE47-0014-01.